OpenStudy (anonymous):

Need help to write the expression as a single logarithm with a coefficient of 1

OpenStudy (anonymous):

\[\log_{10} (x ^{2}-16)-3\log_{10}(x+4)+2\log_{10}x \]

OpenStudy (anonymous):

i need to understand how to do this.. not just the answer.

terenzreignz (terenzreignz):

You need two properties of logs \[\log_{b}M + \log_{b}N=\log_{b}MN\] and \[p \log_{b}M=\log_{b}M^{p}\]

terenzreignz (terenzreignz):

Oh yeah, following from the first property \[\log_{b}M - \log_{b}N=\log_{b}\frac{M}{N}\] provided N is not zero

terenzreignz (terenzreignz):

Actually, provided N > 0

OpenStudy (anonymous):

\[\log_{10}(x ^{2} -16)-\log_{10}(x+3)^{3} +\log_{10}x^2 \]

terenzreignz (terenzreignz):

So far so good. Now combine them all into one log. Hint: If it's a positive log, it goes in the numerator, if it's negative, it goes in the denominator...

OpenStudy (anonymous):

\[\frac{ \log_{10}(x ^{2}-16) }{\log_{10}(x+4)^{3}\times x ^{2} }\] I get confused with the cube and the square. and i probably need to expand the numerator into (x+4)(x-4)??

terenzreignz (terenzreignz):

Wait, let's go over this again... :)

OpenStudy (anonymous):

\[=\log_{10}\frac{ 1 }{(x+4+x ^{2}) } \]?

terenzreignz (terenzreignz):

Once you combine them into one expression, only one log is needed... also, if the log was positive, then it goes in the numerator.

OpenStudy (anonymous):

so x^2 goes in the numerator as well?

terenzreignz (terenzreignz):

Yes it does.

terenzreignz (terenzreignz):

And only one log

OpenStudy (anonymous):

do you expand out the numerator and denominator?

terenzreignz (terenzreignz):

Step by Step: \[\log_{10}(x ^{2} -16)-\log_{10}(x+3)^{3} +\log_{10}x^2\] \[\log_{10}\frac{x^{2}(x^{2}-16)}{(x+3)^{3}}\] this part, you get, right?

terenzreignz (terenzreignz):

And if you don't, don't hesitate to tell me...

OpenStudy (anonymous):

yes i do

terenzreignz (terenzreignz):

Well, too bad, I made a serious typo (and so did you) :P I'm sorry, it should be \[\log_{10}(x ^{2} -16)-\log_{10}(x+4)^{3} +\log_{10}x^2\] \[\log_{10}\frac{x^{2}(x^{2}-16)}{(x+4)^{3}}\]

terenzreignz (terenzreignz):

So, there, from there, you can do your fancy factoring and cancellations :)

terenzreignz (terenzreignz):

What's important is that it's now a single log with coefficient 1,as was required.

OpenStudy (anonymous):

ahh i change it to x+3 heh. ok. \[\log_{10}\frac{ (x-4)x^{2} }{ (x+4)^{2} } \] hmm this isnt right..

terenzreignz (terenzreignz):

Why not?

OpenStudy (anonymous):

it doesnt seem simplified enough... ?

terenzreignz (terenzreignz):

It is.

terenzreignz (terenzreignz):

Unless you prefer to distribute the x^2, and evaluate the square of the binomial at the bottom. Personally, I like to leave it in factored form.

OpenStudy (anonymous):

so that is the final answer???!!

OpenStudy (anonymous):

so if its \[\log_{10}(x ^{2} -4)^{2}-3[\log_{10}(x+4)+2\log_{10}x] \] then i just do whats in the brackts first. so (x+4)(x^2) ?

terenzreignz (terenzreignz):

yes, IF there are brackets.

OpenStudy (anonymous):

ok, and thanks for your help, much appreciated.

terenzreignz (terenzreignz):

No problem :D